OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $70,000. It increased by the same factor each year and in 2006, the average annual salary was $82,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (anonymous):

(A) f(x) = 70(1.17)^x (B) f(x) = 82(1.17)^x (C) f(x) = 70(2.2)^x (D) f(x) = 82(2.2)^x

OpenStudy (anonymous):

Can someone plz help me with this question? @pooja195 @Michele_Laino @iGreen

OpenStudy (anonymous):

You need to find the factor of increase from 2005 to 2006. Since you are given the amounts 70,000 and 82,000 it should be fairly simple to find. this will be you rate or r value.

OpenStudy (anonymous):

820000/70000 = 1.17?

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

P = 82?

OpenStudy (anonymous):

now use that as the value of your (1+r)

OpenStudy (anonymous):

(1 + r) = 2.17?

OpenStudy (anonymous):

No, since the question asks for the rate of increase since 2005, your P should be the amount during 2005

OpenStudy (anonymous):

oh okay so 70

OpenStudy (anonymous):

Yes, and I think 1.17 is the value you want. let me double check .

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So would the answer be A?

OpenStudy (anonymous):

yes use 1.17

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

you're welcome